# Zhuoming Chen Ranajoy Sadhukhan Zihao Ye Yang Zhou Jianyu Zhang Niklas Nolte

### v2

@@ 5.3 Ablation Study
(a) Importance of centering
(b) CWE
(c) FWE
Figure 9: Left: Accuracy comparison for with and without centering. Here we fix $K$ and vary $L$ for the two settings. Mid and Right: Comparison between $\mathrm{TopK}$ attention and MagicPIG. In the two aggregated tasks, sampling-based MagicPIG can even beat the exact $\mathrm{TopK}$ attention. The experiments are done on RULER [33] with a 16K context size.
@@
Centering is important for good performance. In Section 4.3, we use a translation to center the keys before applying LSH sampling. Empirical results show this to be important for downstream tasks as shown in Figure 9a. Without centering, the accuracy drops to almost zero in retrieval (NIAH) and degrades to $65\%$ in FWE. We find almost no keys (less than $0.1\%$) can be sampled by the query without centering, as their orientation is almost opposite, as shown in Figure 2c.
Sampling goes beyond $\mathrm{TopK}$. In Figures 9b and 9c, We compare the performance of MagicPIG and $\mathrm{TopK}$ attention in two aggregated tasks (CWE, FWE) where $\mathrm{TopK}$ attention experiences significant performance degradation (Figure 1). MagicPIG can even beat exact $\mathrm{TopK}$ attention in these two tasks by a margin up to $3\%$ and $8\%$ respectively, demonstrating that sampling improves the ceiling of $\mathrm{TopK}$, which is impossible for a search-only algorithm.
@@ 6 Conclusion
Appendix A Proofs for theorems
A.1 Proof for Theorem 3.2
Proof.
$\mathbb{E}(\bar{o})=\frac{1}{\mathcal{B}}\sum_{j=1}^{\mathcal{B}}\mathbb{E}[v_{i_{j}}]=\frac{1}{\mathcal{B}}\sum_{i=1}^{{n}}w_{i}v_{i}=o$ (12)
Assume $\Sigma_{1}$ is the covariance matrix of $\bar{o}$, $\Sigma_{2}$ is the covariance matrix of $v_{i}$
$\text{Tr}(\Sigma_{1})=\frac{1}{\mathcal{B}}\text{Tr}(\Sigma_{2})=\frac{1}{\mathcal{B}}(\mathbb{E}[||v_{i}||^{2}]-||\mathbb{E}[v_{i}]||^{2})=\frac{1}{\mathcal{B}}(\mathbb{E}[||v_{i}||^{2}]-||o||^{2})$ (13)
$\mathbb{E}[||v_{X}||^{2}]-||o||^{2}$ is a constant, so the trace of covariance matrix monotonically decreases with $\mathcal{B}$. ∎
A.2 Proof for Theorem 3.3
Proof.
$\mathbb{E}[|S|]=\mathbb{E}\Big{[}\sum_{i=1}^{n}\textbf{1}_{i\in S}\Big{]}=\sum_{i=1}^{n}\mathbb{E}[\textbf{1}_{i\in S}]=\sum_{i=1}^{n}(1-(1-w_{i})^{\mathcal{B}})=n-\sum_{i=1}^{n}(1-w_{i})^{\mathcal{B}}$ (14)
Without loss of generality, let $a_{i}=1-w_{i}$ and $a_{1}=\min_{1\leq i\leq n}a_{i}={\epsilon}$, then
$\mathbb{E}[|S|]$ $=n-\sum_{i=1}^{n}a_{i}^{\mathcal{B}}=n-a_{1}^{\mathcal{B}}-\sum_{i=2}^{n}a_{i}^{\mathcal{B}}$ (15)
$=n-{\epsilon}^{\mathcal{B}}-\sum_{i=2}^{n}a_{i}^{\mathcal{B}}$ (16)
$f(x)=x^{\mathcal{B}}$ is convex function with $\mathcal{B}\geq 1$ and $x\geq 0$. Then with Jensen’s inequality, we have
$\sum_{i=2}^{n}a_{i}^{\mathcal{B}}$ $\geq(n-1)\Big{(}\frac{\sum_{i=2}^{n}a_{i}}{n-1}\Big{)}^{\mathcal{B}}=(n-1)\Big{(}\frac{(\sum_{i=1}^{n}a_{i})-a_{1}}{n-1}\Big{)}^{\mathcal{B}}$ (17)
$=(n-1)(\frac{n-1-{\epsilon}}{n-1})^{\mathcal{B}}=(n-1)(1-\frac{{\epsilon}}{n-1})^{\mathcal{B}}$ (18)
Let $g(x)=(1-x)^{\mathcal{B}}+\mathcal{B}x-1$. We can prove $g(x)\geq 0$ for any $x\in(0,1),\mathcal{B}\geq 1$. Then we have
$\sum_{i=2}^{n}a_{i}^{\mathcal{B}}\geq(n-1)(1-\frac{{\epsilon}\mathcal{B}}{n-1})=n-1-{\epsilon}\mathcal{B}$ (19)
Then we finally have
$\mathbb{E}[|S|]=n-{\epsilon}^{\mathcal{B}}-\sum_{i=2}^{n}a_{i}^{\mathcal{B}}\leq 1+{\epsilon}\mathcal{B}$ (20)
Appendix B Oracle sampling
The optimal sampling probability to guarantee estimation is unbiased in terms of lowest variance is not directly using attention score distribution $w_{i}$, but $u_{i}^{\prime}\propto w_{i}||v_{i}||$. However, this sampling probability is not optimal in terms of downstream accuracy and efficiency. We attribute this to two reasons. First, we observe the value norm of the sink token is significantly smaller than others (Figure 11), given its lower probability of being sampled, which may influence the functionality of attention. Second, due to the same reason, $u_{i}^{\prime}\propto w_{i}||v_{i}||$ is flatter than $w_{i}$, resulting larger computation cost (as analyzed by Theorem 3.3).
Appendix C Supplementary analysis
Figure 10: The range of fluctuation of $\log|v_{i}-o|$ and $\frac{qk_{i}^{T}}{\sqrt{d}}$ in a single decoding step. Compared to $\frac{qk_{i}^{T}}{\sqrt{d}}$, $\log|v_{i}-o|$ is stable, hence we do not consider $\log|v_{i}-o|$ in our proposed sampling probability.
Figure 11: The $y$-axis is the norm of values states $\|v_{i}\|$ for token $i$ (on the x-axis). We observe that the value norm $\|v_{0}\|$ of the attention sink is significantly smaller than others.
Figure 10 shows that compared to $\frac{qk_{i}^{T}}{\sqrt{d}}$, $\log|v_{i}-o|$ is stable in a decoding step. Figure 11 shows that the norm of the value states of attention sink is smaller than others.
Appendix D Additional Experiments
D.1 Scaling-up
We provide accuracy results for larger models (meta-llama/Meta-Llama-3.1-70B-Instruct [25]).
Table 4: Synthesized tasks from RULER [33]. MagicPIG preserves high accuracy with low computation for 70B level models. 4 layers {0,16,32,48} are preserved. Config and cost are defined as in Table 1.
Methods Config 16K 32K 64K Avg. $\text{Cost}_{1}$ $\text{Cost}_{2}$ $\text{Cost}_{total}$.
Llama-3.1-70B-Instruct Full 96.4 94.6 88.4 93.1 0.00 1.00 1.00
MagicPIG (10,150) 95.5 93.4 88.2 92.4 0.00 0.02 0.02
D.2 Reasoning
In mathematical reasoning tasks infini_igsm [68, 69], MagicPIG consistently outperforms Quest [58] across all complexity (in terms of operators). We also find $\mathrm{TopK}$ attention suffers from significant performance degradation while Oracle Sampling can maintain high accuracy.
Table 5: Tasks from infini_igsm [68, 69]. MagicPIG preserves high accuracy for reasoning tasks. Config and cost for MagicPIG and Quest are defined as in Table 1. Config denotes the ratio of selected tokens for $\mathrm{TopK}$ and sampled tokens for oracle sampling. For oracle sampling, massive duplication exists in sampled tokens, so $\text{Cost}_{2}$ is significantly lower than the ratio of sampled tokens Theorem 3.3.
Task Methods Config 2-Ops 4-Ops 5-Ops $\text{Cost}_{1}$ $\text{Cost}_{2}$ $\text{Cost}_{total}$.
4K close [68] Llama-3.1-8B-Instruct Full 87.4 71.4 26.8 0.00 1.00 1.00
MagicPIG (10,300) 83.1 67.2 20.7 0.00 0.06 0.06
MagicPIG (10,220) 79.8 58.9 17.9 0.00 0.04 0.04
MagicPIG (10,150) 68.3 43.5 11.7 0.00 0.02 0.02
$\mathrm{TopK}$ 0.06 78.6 62.9 20.8 0.50 0.06 0.56
$\mathrm{TopK}$ 0.04 76.2 59.0 19.2 0.50 0.04 0.54
$\mathrm{TopK}$ 0.02 71.5 44.0 11.3 0.50 0.02 0.52
Oracle Sampling 0.3 88.1 72.4 27.6 0.50 0.02 0.52
Oracle Sampling 0.1 88.5 69.2 26.2 0.50 0.01 0.51
Oracle Sampling 0.02 83.1 57.9 11.9 0.50 0.005 0.505
Quest (16,0.06) 55.8 23.2 5.2 0.06 0.06 0.12
8K close [69] Llama-3.1-8B-Instruct Full 80.2 68.8 26.0 0.00 1.00 1.00
MagicPIG (10,300) 78.6 61.5 25.2 0.00 0.06 0.06
MagicPIG (10,220) 72.2 60.7 20.4 0.00 0.04 0.04
MagicPIG (10,150) 67.1 44.0 11.9 0.00 0.02 0.02
$\mathrm{TopK}$ 0.06 70.2 61.1 22.3 0.50 0.06 0.56
$\mathrm{TopK}$ 0.04 66.9 55.2 20.6 0.50 0.04 0.54
$\mathrm{TopK}$ 0.02 64.7 47.2 15.9 0.50 0.02 0.52
Oracle Sampling 0.3 80.0 67.3 26.2 0.50 0.02 0.52
Oracle Sampling 0.1 76.6 64.1 25.4 0.50 0.01 0.51
Oracle Sampling 0.02 79.0 60.3 20.4 0.50 0.005 0.505
Quest (16,0.06) 54.8 30.0 11.1 0.06 0.06 0.12

### v4

@@ 5.3 Ablation Study
Figure 9: Left: Accuracy comparison for with and without centering. Here we fix $K$ and vary $L$ for the two settings. Mid and Right: Comparison between $\mathrm{TopK}$ attention and MagicPIG. In the two aggregated tasks, sampling-based MagicPIG can even beat the exact $\mathrm{TopK}$ attention. The experiments are done on RULER (Hsieh et al., 2024) with a 16K context size.
@@
Centering is important for good performance. In Section 4.3, we use a translation to center the keys before applying LSH sampling. Empirical results show this to be important for downstream tasks as shown in Figure 9(a). Without centering, the accuracy drops to almost zero in retrieval (NIAH) and degrades to $65\%$ in FWE. We find almost no keys (less than $0.1\%$) can be sampled by the query without centering, as their orientation is almost opposite, as shown in Figure 2(c).
Sampling goes beyond $\mathrm{TopK}$. In Figures 9(b) and 9(c), We compare the performance of MagicPIG and $\mathrm{TopK}$ attention in two aggregated tasks (CWE, FWE) where $\mathrm{TopK}$ attention experiences significant performance degradation (Figure 1). MagicPIG can even beat exact $\mathrm{TopK}$ attention in these two tasks by a margin up to $3\%$ and $8\%$ respectively, demonstrating that sampling improves the ceiling of $\mathrm{TopK}$, which is impossible for a search-only algorithm.